OpenStudy (anonymous):

Change to a fraction or mixed number in simplest form. 66 2/3% = _____

OpenStudy (anonymous):



OpenStudy (anonymous):

\[66\tfrac{2}{3}\%\] is the same as 2/3

OpenStudy (lgbasallote):

change it to improper fraction first...200/3 = 66.67% change to decimal... 0.6667 fraction form = 66/100 = 33/50 don't know if there is a mixed or improper here

OpenStudy (anonymous):

it is \[\frac{3\times 66+2}{3}\%\] \[\frac{200}{3}\%\] and percent means "over 100" so divide by 100 and you get \[\frac{2}{3}\]

OpenStudy (anonymous):

similarly \[33\tfrac{1}{3}\%\] is the same as \[\frac{1}{3}\]which you might have known already

OpenStudy (anonymous):

So 66/2/3 = 2/3%?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

you started with a percent, and ended up with a fraction, not a percent

OpenStudy (anonymous):

1/3 %

OpenStudy (anonymous):

your answer is not a percent, it is a number

OpenStudy (anonymous):

like \[50\%=\frac{1}{2}\]

OpenStudy (anonymous):

or \[75\%=\frac{3}{4}\] etc. your job was to turn the percent into a number

OpenStudy (anonymous):

so just 1/3

OpenStudy (anonymous):

in your case you had \[66\tfrac{2}{3}\%=\frac{2}{3}\]

OpenStudy (anonymous):

i just mentioned that \[33\tfrac{1}{3}\%=\frac{1}{3}\]

OpenStudy (anonymous):

your job is to take a percent and turn it in to a fraction , not turn a percent into some other percent.

OpenStudy (anonymous):

thanks so much so like say 150% I would try to make it into a fraction right?

OpenStudy (anonymous):

-7 1/3 written as a fraction is _____.